# CogGuide: Human-Like Guidance for Zero-Shot Omni-Modal Reasoning

### v2

@@
Abstract—Targeting the issues of “shortcuts” and insufficient contextual understanding in complex cross-modal reasoning of multimodal large models, this paper proposes a zero-shot multimodal reasoning component guided by human-like cognitive strategies centered on an “intent sketch” The component comprises a plug-and-play three-module pipeline—Intent Perceiver, Strategy Generator, and Strategy Selector—that explicitly constructs a “understand–plan–select” cognitive process. By generating and filtering “intent sketch” strategies to guide the final reasoning, it requires no parameter fine-tuning and achieves cross-model transfer solely through in-context engineering. Information-theoretic analysis shows that this process can reduce conditional entropy and improve information utilization efficiency, thereby suppressing unintended shortcut reasoning. Experiments on IntentBench, WorldSense, and Daily-Omni validate the method’s generality and robust gains; compared with their respective baselines, the complete “three-module” scheme yields consistent improvements across different reasoning engines and pipeline combinations, with gains up to approximately +9.51 percentage points, demonstrating the practical value and portability of the “intent sketch” reasoning component in zero-shot scenarios.
@@ INTRODUCTION
Recently, studies have begun to explore multimodal reasoning that simulates human “sketches of thought.” For example, the Sketch-of-Thought framework [17] introduces a human-like cognitive reasoning paradigm that maintains reasoning accuracy while reducing verbose intermediate reasoning; the Machine Mental Imagery approach [18] enables models to incorporate latent visual “imagination” representations during reasoning, simulating the human process of using mental imagery to aid reasoning, thereby improving reasoning efficiency and understanding of complex scenes. Inspired by such research, we propose a human-like cognitive intent-sketch reasoning component that modularly enhances multimodal intent recognition and reasoning performance. The core idea is to decompose and recombine the context of complex problems and to embed human cognitive processes into model reasoning through explicit strategy-planning prompts (i.e., the “intent sketch” mechanism), thereby strengthening the model’s understanding of multimodal information and latent intent. Specifically, we design a plug-and-play three-module pipeline: Intent Perceiver, Strategy Generator, and Strategy Selector. The Intent Perceiver module analyzes video scenes and generates textual intent; the Strategy Generator module produces multiple candidate reasoning strategies based on the context; and the Strategy Selector module evaluates these candidate strategies and selects the optimal one. Finally, a large language model (serving as the reasoning engine) combines the selected strategy with the context and produces an answer by injecting prompts. This component fully leverages the idea of in-context engineering: the entire process requires no parameter fine-tuning of the underlying models; instead, it dynamically injects high-quality strategy prompts to guide the model to focus on the correct lines of thought and cues during reasoning. Under a zero-shot setting, our method can rapidly adapt to different pretrained multimodal models, achieving general performance gains across models. On multiple challenging multimodal reasoning benchmarks (e.g., IntentBench [7], WorldSense[14] and Daily-Omni[15]), zero-shot evaluations have been conducted by prior work [7]. Taking Qwen2.5-Omni as an example, its baseline accuracies on IntentBench, WorldSense, and Daily-Omni are 64.2%, 45.4%, and 47.45%, respectively, serving as a unified comparison baseline. Our method surpasses the original model on all benchmarks (with accuracy improvements of up to 9.51%). Meanwhile, in terms of “effectiveness vs. cost,” compared with training-based methods and without parameter updates or data preparation, we maintain zero training overhead and low latency, resulting in better overall cost and deployment timelines.
The main contributions of this paper are as follows:
  • Human-like cognitive intent-sketch reasoning component: We design a cognitive reasoning component comprising three modules (Intent Perceiver, Strategy Generator, Strategy Selector) that simulates the human cognitive process of “understanding context–planning strategies–selecting solutions.” By explicitly planning and selecting contextual strategies, it enhances reasoning accuracy on complex multimodal tasks. The component is dynamically assembled at inference time, can handle multimodal tasks such as vision and audio, requires no supervision signals or parameter fine-tuning, and effectively improves the model’s understanding of complex scenes and latent intent under zero-shot conditions. Meanwhile, this modular “intent sketch” reasoning component can be embedded as a plug-and-play unit into existing multimodal models; by providing high-quality strategy context, it yields performance gains without fine-tuning, demonstrating cross-model generality and portability.
  • Theoretical analysis from an information-entropy perspective: From an information-theoretic standpoint, we provide a detailed analysis of the working mechanisms of each module, examine how strategy prompts reduce uncertainty in model decisions and improve information utilization efficiency, and present related theoretical derivations, thereby deepening understanding of the principles of context optimization in zero-shot reasoning.
  • Comprehensive multimodal experimental validation: We conduct ablation and comparative experiments on multiple multimodal reasoning benchmarks—including the human-intent understanding task IntentBench and the audio–video collaborative analysis datasets WorldSense and Daily-Omni—to verify that the component can improve the reasoning accuracy of multimodal models without additional training. In addition, by substituting different large reasoning models as the reasoning engine, we validate the component’s generality, highlighting the practicality of strategy prompts in zero-shot reasoning and the value of in-context engineering.
@@
Figure 2. Intent sketch guided zero-shot omni-modal reasoning component: three-module pipeline and information-entropy analysis.
@@
We propose a human-like cognitive intent-sketch reasoning component comprising three serially connected modules: Intent Perceiver, Strategy Generator, and Strategy Selector. The system receives multimodal input $X=(V,A,Q)$, where $V$ denotes a sequence of video frames, $A$ denotes audio, and $Q$ denotes the user’s textual query. The Intent Perceiver compresses $(V,A)$ into a linguistic representation and extracts intent-related cues pertinent to the question; the Strategy Generator, conditioned on the input $\{Q,\text{Intent}\}$, produces $N$ compact and complementary reasoning strategies; the Strategy Selector, given the question (and options, if any), selects from the $N$ candidates the strategy $S^{*}$ that best fits the current question, which then guides the final chain-of-thought reasoning to generate the answer. The entire process mirrors how humans tackle complex tasks—first understanding intent, then planning strategies, and finally selecting the optimal plan—thereby ensuring a deeper grasp of the question and a clearer reasoning trajectory. Ultimately, the selected strategy is used as explicit guidance fused into the model’s reasoning context, thereby influencing and improving the final answer generation.
@@ 2.1 Intent Perceiver: Intent Representation for Strategy Generation
The Intent Perceiver module performs comprehensive analysis and cross-modal evidence fusion for multimodal inputs to distill the intent and key information most closely related to problem solving, thereby clarifying “what the model should plan around.” The module ingests multimodal cues such as video and audio, parses them, identifies the core aspects that the user’s question truly targets, and outputs an intent representation oriented toward downstream reasoning, in order to reduce ambiguity and noise at the question-understanding stage and establish a clear starting point for strategy planning and answer generation.
This module forms an intent representation jointly from the multimodal input and the question: let the input be $X=(V,A,Q)$, where $V$ is video, $A$ is audio, and $Q$ is the question text; the module outputs the intent representation $Z_{IP}=f(X)$. From the perspective of Shannon information theory, $Z_{IP}$ is a task-relevant distillation of $X$, mapping from high-dimensional multimodal perception to a discrete semantic representation. The goal is to provide additional information gain for the subsequent generation of strategy $S$, so that the information on which subsequent reasoning relies is more targeted to the question’s objective. Considering two ways of generating strategies—whether to include the intent representation $Z_{IP}$ (input only $Q$ vs. input $Q,Z_{IP}$)—compared with using only $Q$, adding $Z_{IP}$ is equivalent to providing nonnegative information gain, satisfying:
$MI\left(S;\left(Q,Z_{IP}\right)\right)-MI(S;Q)=MI\left(S;Z_{IP}\mid Q\right)\geq\$0 (1)
From an entropy standpoint, this corresponds to $H(S\mid Q,Z_{IP})\leq H(S\mid Q)$. In practice, we treat $Z_{IP}$ as an information carrier at an intermediate “bottleneck,” embodying the fusion of the question and multimodal cues, so that the statistical relations associated with the strategy become more concentrated, thereby reducing uncertainty in subsequent strategy planning and decision-making.
Overall, the Intent Perceiver helps reduce uncertainty at the problem-understanding stage—that is, it diminishes ambiguity and multiplicity in the question semantics (reducing “conditional entropy”). By focusing on the key information within the multimodal input that is relevant to the answer, the module to some extent removes noise unrelated to the question, increases the association between the extracted cues and the final answer $Y$, and thus lowers the model’s effective uncertainty when facing input $X$. In other words, by clarifying problem intent, the Intent Perceiver module reduces the ambiguity at the starting point of problem solving and, in subsequent reasoning, effectively lowers the initial value of the conditional entropy $H(Y\mid X)$, providing a more determinate starting point for the entire reasoning chain.
2.2 Strategy Generator: Multi-Strategy Generation Based on Semantic Entropy
After obtaining the intent representation $Z_{IP}=f(V,A,Q)$, the Strategy Generator invokes a large language model (LLM) conditioned on $Z_{IP}$ and the question $Q$, configuring it as a “solution-strategy provider” to produce $N$ distinct candidate reasoning strategies $\{S_{1},S_{2},\ldots,S_{N}\}$. These strategies specify only the key lines of reasoning and modes of evidence organization without directly giving the answer, effectively providing comparable “sketches of thought.” This step relies entirely on prompt engineering to activate the pretrained model’s internal knowledge and introduces no additional trainable parameters.
From an information-theoretic perspective, the strategy can be viewed as a random variable $S$. To characterize uncertainty at the “strategy layer,” we define a set of semantic equivalence classes $\mathcal{M}$ over the strategy semantic space, merging strategies that differ in expression but are semantically equivalent into the same class $m\mathcal$$\in$ M}. Given $(Q,Z_{IP})$, the model’s posterior over semantic classes is denoted $p(m\mid Q,Z_{IP})$, from which we define the semantic entropy
[…]
Compared with conditioning on $Q$ alone, explicitly incorporating $Z_{IP}$ amounts to providing additional information such that
$MI(S;Z_{IP}\mid Q)\ =\ H_{sem}(S\mid Q)-H_{sem}(S\mid Q,Z_{IP})\ \geq\ 0$ (3)
Thereby reducing uncertainty at the strategy layer and making the generated strategies more attuned to the video-related context while focusing on the key information needed to solve the current problem.
Generating multiple strategies in one pass can be understood as sampling from the posterior $p(S\mid Q,Z_{IP})$ to cover multiple high–probability-mass reasoning hypotheses. Given the candidate set $\{S_{1},S_{2},\ldots,S_{N}\}$, the downstream step will select the strategy that best fits the current task; even before selection occurs, the presence of multiple candidates already expands the explorable reasoning space and increases the chance of finding a strategy that yields smaller conditional uncertainty $H(Y\mid Q,Z_{IP},S)$. To prevent semantic collapse of candidate strategies onto a single path while maintaining diversity and coverage of the strategy set, we unify single-strategy clarity (low entropy) and set-level coverage (high entropy) within an optimization framework. Specifically, for each candidate $S_{i}$, let its posterior over semantic classes be $p_{i}(m)=p(m\mid Q,Z_{IP};S_{i})$; its single-strategy semantic entropy is then
[…]
@@ 2.2 Strategy Generator: Multi-Strategy Generation Based on Semantic Entropy
where $\alpha,\gamma>0$ are weights, and $Div(\cdot)$ can be defined via pairwise distances in semantic similarity between candidates to encourage complementary strategies with different emphases, such as “evidence-first,” “temporal/causal-first,” or “cross-modal-alignment-first.” Through this process, the Strategy Generator provides a structured, comparable set of candidate lines of thought aligned with $(Q,Z_{IP})$ for subsequent strategy selection and answer generation.
Overall, by explicitly introducing the reasoning-path variable $S$ and sampling multiple possible values for it, the Strategy Generator increases the diversity of candidate strategies and avoids the model’s premature confinement to a single path, enabling a more comprehensive exploration of the solution space. This supplies more comparative information for the subsequent steps and thereby lays the groundwork for reducing $H(Y\mid X)$.
@@
2.3 Strategy Selector: Strategy Selection Based on Minimum Conditional Entropy/Bayesian Risk
[…]
Equivalently, the selection can be viewed as maximizing information gain, from the relationship between conditional entropy and mutual information:
$IG(Y;S_{i}\mid X)=H_{\theta}(Y\mid X)-H_{\theta}(Y\mid X,S_{i})$ (7)
so that:
[…]
Moreover, because the candidates are $N$ complementary strategies generated in the previous stage under $(Q,Z_{IP})$, selection under the above criterion is equivalent to solving over a finite set:
[…]
This mechanism ensures that, when the model ultimately answers the question, it follows a problem-solving approach that is demonstrably more feasible and efficient, thereby increasing the specificity and reliability of reasoning. From a Bayesian chain perspective of the entire pipeline, explicitly conditioning on the “selected strategy” contracts the answer posterior from $p_{\theta}(Y\mid X)$ to $p_{\theta}(Y\mid X,S^{*})$. Combined with the analytical framework that “each additional effective condition reduces uncertainty about the final answer,” the strategy-selection step provides a prior constraint on the reasoning path, further concentrating the posterior and reducing decision ambiguity and empirical error.
@@ 2.4 A Unified Information-Theoretic Framework of Uncertainty Reduction
[…]
@@ 3.1 Experimental Settings
We evaluate the proposed method on multiple multimodal reasoning benchmark datasets, including the human-intent understanding task IntentBench, the audio–video collaborative analysis dataset WorldSense, and the daily-life scenario dataset Daily-Omni. Accuracy (%) is used as the evaluation metric to compare the baseline performance of the original models (without any strategy prompts) against the reasoning performance after introducing different combinations of modules. The method consists of three pluggable modules—Intent Perceiver, Strategy Generator, and Strategy Selector—and the final output is fed into a reasoning engine to produce answers (see Table 1).
Each experiment is conducted under a zero-shot setting, i.e., no additional fine-tuning of the models and zero-shot prompts are used. We select three large multimodal models as reasoning engines: HumanOmniV2, Qwen2.5-Omni, and Qwen2.5-VL, which represent different categories of multimodal large models. Our method employs four pretrained language models as pipeline module models (Pipeline): the closed-source GPT-4o and Doubao-Seed-1.6, and the open-source GLM-4.5 and Qwen3.
To quantify the contribution of each front-end module to the final reasoning performance, we conduct ablation studies centered on the three modules—Intent Perceiver, Strategy Generator, and Strategy Selector (see Table 2). CG_Qwen_vl indicates the complete three-module configuration: take Qwen2.5-VL-32B as intent perceiver, then three candidate strategies are produced, and the selector is enabled to choose the best; CG_GLM_vl take GLM-4.5V as the intent perceiver model to test the substitutability of the intent perceiver; Abl_No_Intent disables the intent perceiver module and inputs the question text directly into strategy generation; Abl_Single_Strategy sets the number of strategies to 1 and disables the selector, thereby isolating the gain from “multiple strategies + selection”; BaseLine completely removes the front-end pipeline (no intent perceiver, no strategy generation, no selection), and the reasoning engine answers solely based on the original multimodal input and the question. Each group is evaluated under the Cartesian product of the four Pipeline models (GPT-4o / GLM-4.5 / Doubao-Seed-1.6 / Qwen3) and the three reasoning engines (HumanOmniV2 / Qwen2.5-Omni / Qwen2.5-VL), with other hyperparameters and inference configurations kept consistent to ensure comparability. Through these various combinations, we verify the magnitude of performance improvement when all three modules are integrated and assess the contribution of each module to overall performance.
@@
GPT-4o Strategy Generator/Strategy Selector Large Closed-Source Model
GLM-4.5 Strategy Generator/Strategy Selector 355B/32B
Doubao-Seed-1.6 Strategy Generator/Strategy Selector Large Closed-Source Model
Qwen3 Strategy Generator/Strategy Selector 235B/22B
@@
Table2. Experimental Configurations: Three-Module Settings and Descriptions
Experiment ID Intent Perceiver Strategy Generation Strategy Selection Description
CG_Qwen_vl Qwen2.5-VL-32B 3 On All three modules enabled; use Qwen as Intent model
CG_GLM_vl GLM-4.5V 3 On All three modules enabled; use GLM as Intent model
Abl_NoIntent No 3 On Remove the Intent module
Abl_SingleStrategy Qwen2.5-VL-32B 1 No Change strategy generation to single strategy
BaseLine No No No No front-end pipeline

### v3

@@
Abstract—Addressing shortcut reliance and limited contextual understanding in cross-modal reasoning, we introduce a zero-shot omni-modal reasoning component inspired by human-like cognition, realized as a plug-and-play intent-sketch pipeline with three serial modules: Intent Perceiver, Policy Generator and Strategy Selector, that models an understand-plan-select cognitive process. By producing and filtering lightweight intent sketch strategies to guide reasoning, the method requires no parameter fine-tuning and enables cross-model transfer through in-context engineering. An information-theoretic analysis demonstrates that this process reduces conditional entropy and improves information utilization efficiency. Extensive experiments on IntentBench, WorldSense, and Daily-Omni confirm the method’s generality and robustness: the full three-module design consistently outperforms strong baselines across diverse reasoning engines and pipeline settings, with maximum gains of +9.51 pp and a relative improvement of 20.04%. These results demonstrate the practical value and portability of the proposed intent sketch reasoning component for zero-shot omni-modal reasoning.
@@ INTRODUCTION
Recently, studies have begun to explore multimodal reasoning that simulates human “sketches of thought.” For example, the Sketch-of-Thought framework [17] introduces a human-like cognitive reasoning paradigm that maintains reasoning accuracy while reducing verbose intermediate reasoning; the Machine Mental Imagery approach [18] enables models to incorporate latent visual “imagination” representations during reasoning, simulating the human process of using mental imagery to aid reasoning, thereby improving reasoning efficiency and understanding of complex scenes. Inspired by such research, we propose a human-like cognitive intent-sketch reasoning component that modularly enhances multimodal intent recognition and reasoning performance. The core idea is to decompose and recombine the context of complex problems and to embed human cognitive processes into model reasoning through explicit strategy-planning prompts (i.e., the “intent sketch” mechanism), thereby strengthening the model’s understanding of multimodal information and latent intent. Specifically, we propose a zero-shot omni-modal reasoning component inspired by human-like cognition. The component decomposes context through a plug-and-play intent-sketch pipeline: Intent Perceiver, Policy Generator and Strategy Selector, which infer textual intent from video and audio, generate candidate policies, and select an optimal strategy; a reasoning LLM then conditions on the selected strategy to produce the final answer via prompt composition. On multiple challenging multimodal reasoning benchmarks (e.g., IntentBench [7], WorldSense[14] and Daily-Omni[15]), using HumanOmniV2, Qwen2.5-Omni, Qwen2.5-VL, respectively, serving as a unified comparison. Our method surpasses the base model on all benchmarks (highest relative increase: 20.04%). Meanwhile, in terms of “effectiveness vs. cost” compared with training-based methods, we maintain zero training overhead and low latency, resulting in better overall cost and deployment timelines.
The main contributions: (1) a zero-shot omni-modal reasoning component that contains a plug-and-play intent-sketch pipeline, deliver immediate accuracy gains and cross-model generality; (2) an information-entropy analysis that explains how strategy prompts reduce decision uncertainty; (3) present comprehensive omni-modal experiments, including ablations and reasoning-engine swaps, that validate the effectiveness and practicality of in-context strategy prompting on the three benchmarks above.
@@
Figure 2. Zero-shot omni-modal reasoning component: plug-and-play intent–sketch pipeline and information-entropy analysis
@@
The plug-and-play intent–sketch pipeline contains three serial modules: Intent Perceiver, Policy Generator, and Strategy Selector, operating on omni-modal input $X=(V,A,Q)$ where $V$ is video, $A$ is audio, and $Q$ is the textual query. The Intent Perceiver maps $X$ to $Z_{\mathrm{IP}}$ to reduce planning uncertainty; the Policy Generator yields $N$ semantically diverse policies (strategy sketches); the Strategy Selector applies extra conditioning $C$ and selects $S^{*}$ to drive the final reasoning. This design mirrors the human routine of understand–plan–select, yielding a clearer reasoning trajectory and improved answer generation.
@@ 2.1 Intent Perceiver: Intent Representation for Strategy Generation
Given $X=(V,A,Q)$, the module outputs an intent representation $Z_{\mathrm{IP}}=f(X)$ that fuses cross-modal cues with the query. From an information-theoretic view, using $Z_{\mathrm{IP}}$ to condition strategy generation provides nonnegative information gain:
$\mathrm{MI}\!\left(S;\,(Q,Z_{\mathrm{IP}})\right)-\mathrm{MI}(S;Q)=\mathrm{MI}\!\left(S;Z_{\mathrm{IP}}\mid Q\right)\geq 0,$ (1)
which is equivalent to $H(S\mid Q,Z_{\mathrm{IP}})\leq H(S\mid Q)$. By focusing on answer-relevant omni-modal evidence, $Z_{\mathrm{IP}}$ reduces ambiguity at problem understanding and effectively lowers the initial conditional uncertainty $H(Y\mid X)$.
2.2 Policy Generator: Omni-Policy Generation Based on Semantic Entropy
Conditioned on $(Q,Z_{\mathrm{IP}})$, a frozen LLM (prompted as a policy provider) generates $N$ candidate sketches $\{S_{1},\ldots,S_{N}\}$ that specify lines of reasoning without producing the answer. Let $\mathcal{M}$ denote semantic equivalence classes over policies and $p(m\mid Q,Z_{\mathrm{IP}})$ the posterior over classes. We define the semantic entropy
$H_{\mathrm{sem}}(S\mid Q,Z_{\mathrm{IP}})=-\sum_{m\in\mathcal{M}}p(m\mid Q,Z_{\mathrm{IP}})\,\log p(m\mid Q,Z_{\mathrm{IP}}),$ (2)
and the gain brought by intent conditioning:
$\mathrm{MI}(S;Z_{\mathrm{IP}}\mid Q)=H_{\mathrm{sem}}(S\mid Q)-H_{\mathrm{sem}}(S\mid Q,Z_{\mathrm{IP}})\ \geq\ 0.$ (3)
To balance single-sketch clarity and set-level coverage, for each candidate $S_{i}$ let $p_{i}(m)=p(m\mid Q,Z_{\mathrm{IP}};S_{i})$ and define
$H_{i}=-\sum_{m}p_{i}(m)\,\log p_{i}(m).$ (4)
For the mixture $\bar{p}(m)=\tfrac{1}{N}\sum_{i=1}^{N}p_{i}(m)$ with entropy $H(\bar{p})=-\sum_{m}\bar{p}(m)\log\bar{p}(m)$, we optimize
$\max_{S_{1},\ldots,S_{N}}\ \ H(\bar{p})\;-\;\alpha\,\frac{1}{N}\sum_{i=1}^{N}H_{\mathrm{sem}}(S_{i}\mid Q,Z_{\mathrm{IP}})\;+\;\gamma\,\mathrm{Div}(S_{1},\ldots,S_{N}),$ (5)
@@ 2.2 Policy Generator: Omni-Policy Generation Based on Semantic Entropy
where $\alpha,\gamma>0$ are weights, and $\mathrm{Div}(\cdot)$ can be defined via pairwise distances in semantic similarity between candidates to encourage complementary policies with different emphases, such as “evidence-first,” “temporal/causal-first,” or “cross-modal-alignment-first.”
@@
2.3 Strategy Selector: Selection via Minimum Conditional Entropy / Bayesian Risk
Given $\{S_{1},\ldots,S_{N}\}$, the model (prompted as a strategy evaluator) selects $S^{*}$ based on the input $X$ and task-specific conditioning $C$ to align the choice with real-world constraints. Let $p_{\theta}(Y\mid X,S_{i},C)$ be the answer posterior under $S_{i}$ with extra conditioning $C$. In implementation, we obtain $p_{\theta}(Y\mid X,S_{i},C)$ by prompting the LLM with $(X,S_{i},C)$ and computing answer-slot likelihoods, so the added $C$ induces strategy-dependent posteriors. Under 010\!-\!1 loss,
$S^{*}=\arg\min_{i}R_{0\text{--}1}(S_{i})=\arg\max_{i}\ \max_{Y}\ p_{\theta}(Y\mid X,S_{i},C).$ (6)
Selection maximizes information gain:
$IG(Y;S_{i}\mid X)=H_{\theta}(Y\mid X)-H_{\theta}(Y\mid X,S_{i},C),$ (7)
thus
$S^{*}=\arg\max_{i}\ IG(Y;S_{i}\mid X)=\arg\min_{i}\ H_{\theta}(Y\mid X,S_{i},C).$ (8)
Because candidates are complementary,
$H_{\theta}(Y\mid X,S^{*})=\min_{i}\ H_{\theta}(Y\mid X,S_{i},C)\ \leq\ \frac{1}{N}\sum_{i=1}^{N}H_{\theta}(Y\mid X,S_{i},C).$ (9)
In practice, exploring a diverse and complementary set of strategies increases the chance that at least one $S_{i}$ substantially reduces uncertainty, making lower overall uncertainty more likely. This implies a lower expected conditional entropy—and, via a Fano-type bound, a lower achievable error bound; in particular, reducing $H_{\theta}(Y\mid X,S^{*},C)$ lowers an upper bound on the minimum error rate under 010\!-\!1 loss. Consequently, subsequent reasoning based on $S^{*}$ integrates omni-modal evidence along a path of higher confidence, enhancing the reliability of answer generation.
@@ 2.4 A Unified Information-Theoretic Framework of Uncertainty Reduction
This section formalizes our pipeline as an information-theoretic mechanism for uncertainty reduction. We begin with the basic theorem of conditioning:
$H(X\mid Y)\ \leq\ H(X).$ (10)
Treating the selected strategy $S^{*}$ as an observable planning variable, and letting $C$ denote the extra task conditioning used to choose $S^{*}$, we model the answer posterior as $p(Y\mid X,I,S,C)$, where $I\equiv Z_{\mathrm{IP}}$. By the conditioning-reduces-entropy theorem, we obtain the following monotone contraction chain:
$H(Y\mid X)\ \geq\ H(Y\mid X,I)\ \geq\ H(Y\mid X,I,S)\ \geq\ H(Y\mid X,I,S,C).$ (11)
This captures how intent $I$, the multi-policy set $S$, and the selected strategy $S^{*}$ progressively reduce the conditional uncertainty of $Y$.
From the perspective of the Data Processing Inequality (DPI), if $X\!\to\!I\!\to\!S\!\to\!S^{*}$ forms a Markov chain, then
$I(X;S^{*})\ \leq\ I(X;S)\ \leq\ I(X;I)\ \leq\ I(X;X),$ (12)
which implies that any representation derived from $X$ must lose some information about $X$ itself. However, in inference, we do not replace $X$ with $S^{*}$; instead, we use $S^{*}$ as auxiliary information conditioned on $X$:
$H(Y\mid X,S^{*})\ <\ H(Y\mid X),$ (13)
with strict inequality when $I(Y;S^{*}\mid X)>0$. This reduction occurs because $S^{*}$, although derived from $X$, is constructed through a process that incorporates task-specific guidance (e.g., evidence-first, causal-first, cross-modal alignment). These human-informed design choices embed additional relevant cues into $S^{*}$, enabling it to provide new information about $Y$ that is not directly accessible from $X$ alone. Thus, conditioning on $S^{*}$ enhances prediction accuracy, leading to the entropy reduction in (13) and contributing to the overall uncertainty contraction in (11).
@@ 3.1 Experimental Settings
We evaluate the method on three omni-modal reasoning benchmarks: IntentBench (omni-intent understanding), WorldSense (audio-video collaborative analysis), and Daily-Omni (daily-life scenarios). Accuracy (%) is the metric. We compare baseline models (no strategy prompts) with our approach that adds three plug-in modules—Intent Perceiver, Policy Generator, and Strategy Selector—whose output is passed to the reasoning engine (Table I).
All experiments follow a zero-shot setting (no fine-tuning; zero-shot prompts only). We use three reasoning engines: HumanOmniV2, Qwen2.5-Omni, and Qwen2.5-VL. The pipeline modules are instantiated with four pretrained LLMs: closed-source GPT-4o and Doubao-Seed-1.6, and open-source GLM-4.5 and Qwen3. To quantify module contributions, we conduct ablations (Table II). Each setting is evaluated over the Cartesian product of the four pipeline LMs and the three reasoning engines, with all other hyperparameters and inference configurations held constant for fair comparison. Through these combinations we measure the overall gain of the integrated system and the contribution of each module.
@@
GPT-4o Policy Generator/Strategy Selector Large Closed-Source Model
GLM-4.5 Policy Generator/Strategy Selector 355B/32B
Doubao-Seed-1.6 Policy Generator/Strategy Selector Large Closed-Source Model
Qwen3 Policy Generator/Strategy Selector 235B/22B
@@
Table 2. Experimental Configurations: Three-Module Settings
Experiment ID Intent Perceiver Policy Generation Strategy Selection
CG_Qwen Qwen2.5-VL-32B 3 On
CG_GLM GLM-4.5V 3 On
Abl_NI No 3 On
Abl_SP Qwen2.5-VL-32B 1 On
BaseLine No No No
Notes. CG_Qwen: Full pipeline, use Qwen as Intent model. CG_GLM: Full pipeline, use GLM as Intent model. Abl_NI: Remove the Intent module. Abl_SP: Change policy generation to single policy. BaseLine: No front-end pipeline.